OpenStudy (anonymous):

{} = {x|x is a real number and x^2=-1} True or false? why?

OpenStudy (raden):

false.. if x a real number (maybe positive, zero or a negative number) then x^2 must be a postive number and impossible to get a negative number

OpenStudy (raden):

in quadratic function : y = ax^2+bx+c for the value the discriminant (D) < 0, its called definite positive with D=b^2-4ac so, x^2 = -1 or x^2 + 1 = 0 D=0^2-4(1)(1) = -4

OpenStudy (unklerhaukus):

\[x=\pm i, \qquad \pm i\not\in \mathbb R\]

OpenStudy (anonymous):

haha~ true..coz x^2 cannot be -1..thxs a lot..

OpenStudy (raden):

yepss.. :)